OpenStudy (anonymous):

Evaluate 15 - 4 + (7 - 5)^2 = __________.

OpenStudy (anonymous):

use BEDMAS.

OpenStudy (anonymous):

=15 - 4 + (2)^2 = 15 - 4 + 4 = 15

OpenStudy (hockeychick23):

ok so again you use PEMDAS for this so we focus on the parentheses first. So we know 7-5=2 so now we need to do 2^2 or 2*2 which is 4 so now you have: 15-4+4

OpenStudy (anonymous):

so we be back in 15 after 15-4 plus 4

OpenStudy (hockeychick23):

yep we have 15-4+4. and -4+4= 0 so we know that the answer is 15